OpenStudy (anonymous):

-6(+-5)-(++5)=6 solve andget a solution

OpenStudy (anonymous):

Step 1: simplify what's inside the brackets. + multiplied by + = + + multiplied by - = - - multiplied by - = Can you resolve that?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

okay lets look at the first bracket You have (+-) that is + multiplied by - right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Looking at my first post where i gave you the laws , what is the resilt of +-

OpenStudy (anonymous):

-

OpenStudy (anonymous):

correct ! so the term in brackets can be written as (-5) What about the second term, refer to the multiplication laws again.

OpenStudy (anonymous):

- -

OpenStudy (anonymous):

in the second term , inside the brackets you have (++5) can you see that/ what is + multiplied by + ?

OpenStudy (anonymous):

+

OpenStudy (anonymous):

great...So in the equation is -6(-5)-(+5)=6

OpenStudy (anonymous):

or simply -6(-5)-(5)=6

OpenStudy (anonymous):

Now you multiply the -6(-5)...what do you get?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

30 the answer

OpenStudy (anonymous):

now 30 -(5) = 25?

OpenStudy (anonymous):

This is an inequality...is your question correct?

OpenStudy (anonymous):

I do not know I do t have tt option

OpenStudy (anonymous):

I am not sure now...do you want to get help elsewhere. Sorry

OpenStudy (anonymous):

I am tired and been worki a y I will try tomorw because they are shutting down open study

OpenStudy (anonymous):

When?

OpenStudy (anonymous):

at 9pm

OpenStudy (anonymous):

its 1am here in ireland how many minutes befor 9pm

OpenStudy (anonymous):

oh wow they said in a hour

OpenStudy (anonymous):

good luck..if i find a solution I will post it here

OpenStudy (anonymous):

or at cdelomas@yahoo.com

OpenStudy (anonymous):

okay...keep well

OpenStudy (anonymous):

thanks